OpenStudy (shelby1290):

5x=4x+7 would a short cut to doing this question be 7-5 divided by 4 ?

OpenStudy (the_fizicx99):

No?

OpenStudy (shelby1290):

so would it be the manuel way (long way)

OpenStudy (shelby1290):

that involves dividing both sides

OpenStudy (anonymous):

no. there is no long way this only takes 1 step to solve

OpenStudy (ganpat):

u can try that.. Just u have messed up the numbers.. Try and Re arrange !!

OpenStudy (shelby1290):

5x-4x

OpenStudy (shelby1290):

that answer divided by 7

OpenStudy (shelby1290):

sorry i just really do suck at this unit

OpenStudy (ganpat):

perfect !!.. thats solution for x !!

OpenStudy (anonymous):

why are you dead set on dividing by 7 just use inverse operations on the 4x across the = sign and you have your answer

OpenStudy (shelby1290):

dont you have to divide by 7 ?

OpenStudy (shelby1290):

but isnt that what i just did

OpenStudy (ganpat):

oops.. i read it wrong.. divide the 7 by answer.. as @lonnie455rich said..

OpenStudy (shelby1290):

so the answer is 7 divided by 5 ?

OpenStudy (shelby1290):

x=1.4?

OpenStudy (ganpat):

No !! 5x = 4x + 7 5x - 4x = 7 1x = 7 x = 7/1 therefore, x = 7

OpenStudy (ganpat):

can u just tell me, how did u get x = 1.4 ?? M just Curious.. :) :D

OpenStudy (anonymous):

remember the 1 is understood by x so we just don't even write it so its just x=7 after you subtract the 4x from both sides using inveroperations

OpenStudy (shelby1290):

ohhh okay wait what if i had to do 11s+25=54+25+2s

OpenStudy (anonymous):

inverse operations that is all you need to know when solving MOST types of equations

OpenStudy (shelby1290):

7 divided by 5

OpenStudy (shelby1290):

sorry i did some extreme weirdness

OpenStudy (shelby1290):

im just not so great at math

OpenStudy (ganpat):

put all s terms on one side and rest all terms on other side.. (Try solving)

OpenStudy (anonymous):

first thing I would do on that one is examine the 25s. then I would get your variable on 1 side then I would use inverse operations to finish solving

OpenStudy (ganpat):

hey relax !! m just kidding man.. u doing good.. u need to practise a little :)

OpenStudy (shelby1290):

@lonnie455rich by inverse you mean opposite ? right

OpenStudy (anonymous):

yes like + - / * radical powers exponential log derivative integral and so on

OpenStudy (shelby1290):

ahh okay and collect like terms? or just divide by 11

OpenStudy (anonymous):

don't divide by 11 until you get all your variables together. and get your contants all on the same side

OpenStudy (shelby1290):

so ... 11s+2s=25-54+25

OpenStudy (shelby1290):

^@Ganpat

OpenStudy (anonymous):

wow. you messed up signs or something here